# Discrete Time Credit-Based Shaping for Time-Sensitive Applications in 5G/6G Networks

### v2

@@ I Introduction
Beyond-5G (B5G) and 6G networks are expected to support a wide range of quality-of-service (QoS) requirements [Vision6G]. Emerging Industry $4.0/5.0$ use cases such as smart factories and real-time motion control require end-to-end latencies below $1~\text{ms}$ with microsecond-level jitter, and reliabilities exceeding $99.999\%$ [Industry5O]. On wired links, the IEEE Time-Sensitive Networking (TSN) toolset provides such determinism by combining per-flow shaping with time-aware queuing to yield hard delay bounds [Sasiain]. Achieving comparable guarantees over the stochastic wireless medium is challenging due to time-varying channel quality, grant-based scheduling, and retransmissions [Sharma]. A natural approach is to adapt TSN primitives to 5G/6G wireless networks, notably the IEEE 802.1Qav [standard] credit-based shaping (CBS), which achieves per-class bounded delay when the aggregate reserved rate is strictly below the link capacity [fi16090345]. This will extend predictable, bounded latency to high-priority traffic, while preventing starvation of lower-priority traffic.
TSN-over-5G deployments introduce device-side and network-side TSN translators (DS-TT/NW-TT) that map IEEE 802.1Q semantics onto 5G QoS identifier (5QI) [Nikhileswar, Zhang, Atiq]. Complementary efforts build TSN–5G testbeds and virtualized environments [garbugli1, garbugli12]. These systems enable end-to-end TSN connectivity but treat the 5G cell as a transparent bridge [Ana2] and leave the NR medium access control (MAC) scheduler unmodified. Closer to the scheduler, configured-grant (CG) approaches coordinate 5G grants with TSN flow periodicity to increase the fraction of flows meeting latency targets [Ana], coupling TSN knowledge to grant timing. Wireless TSN surveys flag slot-level scheduling under variable MCS as an open problem [SurveyWirelessTSN2025, Zanbouri]. Within ultra-reliable low-latency communications (URLLC), mechanisms such as grant-free access, mini-slots, and HARQ tuning improve latency but can starve low-priority flows when critical flows are over-provisioned [URRLCsurvey2025, MaghsoudniaHotNets2024]. Even with DS-TT/NW-TT that map Ethernet priorities into 5QI, the gNB MAC continues to run the standard, non-credit-aware grant scheduler [Satka]. As a result, CBS-derived determinism can be lost inside the NR air interface, where bursts accumulate and slot-level pacing is not enforced. What is missing is a slot-native credit mechanism inside the gNB that debits credit whenever a new transmission grant is issued. We address this gap with a credit gate embedded in the gNB: per-UE credit evolves in discrete time, debits on each issued grant, and can be tied either to granted transport block size (TBS) or to the actually delivered bytes (to avoid over-debiting under partial usage). This enables TSN-like latency isolation in NR while preserving spectral efficiency. Our main contributions can be summarized as follows:
@@ II Background: Credit-Based Shaping in Time-Sensitive Networking
The CBS [standard, Bril] offers per-class bandwidth reservation and contributes to predictable latency by ensuring bounded queuing delays on shared links, working in conjunction with other mechanisms like the time-aware shaper (TAS) for tight delay bounds in Ethernet. Each prioritized traffic class is assigned a dedicated queue, indexed by $p$. Each queue maintains a continuous-time credit, $C_{p}(t)$, and packets are transmitted only when $C_{p}(t){\geq}0$ [Hassani]. Among queues that are both backlogged (i.e., have data to send) and meet this eligibility condition of having non-negative credit, CBS strictly prioritizes the highest priority queue for transmission, allowing preferential treatment and deterministic latency control for critical traffic.
@@ II-A Credit Dynamics: IdleSlope and SendSlope
If queue $p$ is idle (either no pending load and $C_{p}(t)\geq 0$, or $C_{p}(t)<0$ and not transmitting, or pending load but blocked by other traffic), its credit increases. For any two time instances $t_{1},t_{2}$ and $t_{2}{>}t_{1}$ the credit updates as [Jingyue]:
@@
The credit is reset to zero if the queue becomes empty and $C_{p}(t)\geq 0$ (Eq. 3.18 in [Jingyue]).
@@ V Credit-Based Shaping with Partial Usage
A key challenge arises with the CBS-DT adaptation due to the practicalities of 5G NR scheduling. While the gNB knows the precise downlink queue size, the selected $\text{TBS}[n]$ must conform to the quantized values mandated by 3GPP [3gpp38214]. Consequently whenever the granted $\text{TBS}_{u}[n]$ exceeds $Q_{u}[n]$, padding is added. CBS-DT, by debiting the entire grant $\text{TBS}_{u}[n]$ (10), therefore over-penalizes UE credit. This drives the credit deeper into deficit postponing the next eligibility event, delaying subsequent transmission eligibility particularly affecting UEs with bursty arrival traffic patterns. We therefore introduce Credit-Based Shaping with Partial Usage (CBS-PU), retaining the CBS-DT structure outlined in Section IV. The sole modification lies in the definition of the debit term $D_{u}[n]$ used in (7).
@@ VII Performance Evaluation
To evaluate the performance of the proposed CBS-DT and CBS-PU mechanisms, we built a simulation framework upon the MATLAB 5G Toolbox. We simulate a single-cell FDD NR downlink with a 15 kHz subcarrier spacing (SCS), using the 3GPP Urban Macro (UMa) propagation model (TR 38.901); CQI reports are mapped to MCS according to 3GPP TS 38.214. Six static UEs are evenly split across three priority classes: $p_{1}$ (high), $p_{2}$ (medium), and $p_{3}$ (low). Each UE generates a packet stream with nominal rate of $\lambda=450$ pkt/s and a fixed payload size $S_{u}\in\{80,160,240\}$ bytes for $p_{1}$, $p_{2}$, and $p_{3}$, respectively. The aggregate offered rate is defined as $\Lambda_{\text{DL}}=\sum_{u}\lambda S_{u},$ and we report the normalized load $\rho=\Lambda_{\text{DL}}/C_{\text{DL}}$; we study both nominal load ($\rho=1.0$) and a controlled overload ($\rho=4.0$) to stress-test the scheduler’s priority protection capabilities. Class reservations (idleSlope shares) are set to {75%, 20%, 5%} for priorities $p_{1},p_{2},p_{3}$}, respectively. Per-UE credit clamps $lo_{u},hi_{u}]$ are derived following the Ethernet CBS standard [standard]. The credit gate renders a UE eligible in slot $n$ iff it is backlogged, has a free HARQ process, and its credit is non-negative. HARQ retransmissions consume radio resources but do not reduce the downlink queue $Q_{u}$; therefore, only the bytes from a first transmission debit the credit. The CBS-PU mechanism relies on information readily available at the gNB. At the start of each slot, the gNB knows the queue size $Q_{u}[n]$, the chosen TBS, and the HARQ state. A new transmission service removes $\min\{\text{TBS}_{u}[n],Q_{u}[n]\}$ bytes from the buffer. Any padding added to the TB does not change $Q_{u}$.
@@ Acknowledgement
This work was funded by the Dutch ”Rijksdienst voor Ondernemend Nederland – RVO” under the TSH number 21007 – RHIADA project.

### v3

@@ I Introduction
Beyond-5G (B5G) and 6G networks are expected to support a wide range of quality-of-service (QoS) requirements [15]. Emerging Industry $4.0/5.0$ use cases such as smart factories and real-time motion control require end-to-end latencies below $1~\text{ms}$ with microsecond-level jitter, and reliabilities exceeding $99.999\%$ [19]. On wired links, the IEEE Time-Sensitive Networking (TSN) toolset provides such determinism by combining per-flow shaping with time-aware queuing to yield hard delay bounds [16]. Achieving comparable guarantees over the stochastic wireless medium is challenging due to time-varying channel quality, grant-based scheduling, and retransmissions [18]. A natural approach is to adapt TSN primitives to 5G/6G wireless networks, notably the IEEE 802.1Qav [9] credit-based shaping (CBS), which achieves per-class bounded delay when the aggregate reserved rate is strictly below the link capacity [10]. This will extend predictable, bounded latency to high-priority traffic, while preventing starvation of lower-priority traffic.
TSN-over-5G deployments introduce device-side and network-side TSN translators (DS-TT/NW-TT) that map IEEE 802.1Q semantics onto 5G QoS identifier (5QI) [14, 2]. Complementary efforts build TSN–5G testbeds and virtualized environments [6, 5]. These systems enable end-to-end TSN connectivity but treat the 5G cell as a transparent bridge [11] and leave the NR medium access control (MAC) scheduler unmodified. Closer to the scheduler, configured-grant (CG) approaches coordinate 5G grants with TSN flow periodicity to increase the fraction of flows meeting latency targets [12], coupling TSN knowledge to grant timing. Wireless TSN surveys flag slot-level scheduling under variable MCS as an open problem [20]. Within ultra-reliable low-latency communications (URLLC), mechanisms such as grant-free access, mini-slots, and HARQ tuning improve latency but can starve low-priority flows when critical flows are over-provisioned [7, 13]. Even with DS-TT/NW-TT that map Ethernet priorities into 5QI, the gNB MAC continues to run the standard, non-credit-aware grant scheduler [17]. As a result, CBS-derived determinism can be lost inside the NR air interface, where bursts accumulate and slot-level pacing is not enforced. What is missing is a slot-native credit mechanism inside the gNB that debits credit whenever a new transmission grant is issued. We address this gap with a credit gate embedded in the gNB: per-UE credit evolves in discrete time, debits on each issued grant, and can be tied either to granted transport block size (TBS) or to the actually delivered bytes (to avoid over-debiting under partial usage). This enables TSN-like latency isolation in NR while preserving spectral efficiency. Our main contributions can be summarized as follows:
@@ II Background: Credit-Based Shaping in Time-Sensitive Networking
The CBS [9, 3] offers per-class bandwidth reservation and contributes to predictable latency by ensuring bounded queuing delays on shared links, working in conjunction with other mechanisms like the time-aware shaper (TAS) for tight delay bounds in Ethernet. Each prioritized traffic class is assigned a dedicated queue, indexed by $p$. Each queue maintains a continuous-time credit, $C_{p}(t)$, and packets are transmitted only when $C_{p}(t){\geq}0$ [8]. Among queues that are both backlogged (i.e., have data to send) and meet this eligibility condition of having non-negative credit, CBS strictly prioritizes the highest priority queue for transmission, allowing preferential treatment and deterministic latency control for critical traffic.
@@ II-A Credit Dynamics: IdleSlope and SendSlope
If queue $p$ is idle (either no pending load and $C_{p}(t)\geq 0$, or $C_{p}(t)<0$ and not transmitting, or pending load but blocked by other traffic), its credit increases. For any two time instances $t_{1},t_{2}$ and $t_{2}{>}t_{1}$ the credit updates as [4]:
@@
The credit is reset to zero if the queue becomes empty and $C_{p}(t)\geq 0$ (Eq. 3.18 in [4]).
@@ V Credit-Based Shaping with Partial Usage
A key challenge arises with the CBS-DT adaptation due to the practicalities of 5G NR scheduling. While the gNB knows the precise downlink queue size, the selected $\text{TBS}[n]$ must conform to the quantized values mandated by 3GPP [1]. Consequently whenever the granted $\text{TBS}_{u}[n]$ exceeds $Q_{u}[n]$, padding is added. CBS-DT, by debiting the entire grant $\text{TBS}_{u}[n]$ (10), therefore over-penalizes UE credit. This drives the credit deeper into deficit postponing the next eligibility event, delaying subsequent transmission eligibility particularly affecting UEs with bursty arrival traffic patterns. We therefore introduce Credit-Based Shaping with Partial Usage (CBS-PU), retaining the CBS-DT structure outlined in Section IV. The sole modification lies in the definition of the debit term $D_{u}[n]$ used in (7).
@@ VII Performance Evaluation
To evaluate the performance of the proposed CBS-DT and CBS-PU mechanisms, we built a simulation framework upon the MATLAB 5G Toolbox. We simulate a single-cell FDD NR downlink with a 15 kHz subcarrier spacing (SCS), using the 3GPP Urban Macro (UMa) propagation model (TR 38.901); CQI reports are mapped to MCS according to 3GPP TS 38.214. Six static UEs are evenly split across three priority classes: $p_{1}$ (high), $p_{2}$ (medium), and $p_{3}$ (low). Each UE generates a packet stream with nominal rate of $\lambda=450$ pkt/s and a fixed payload size $S_{u}\in\{80,160,240\}$ bytes for $p_{1}$, $p_{2}$, and $p_{3}$, respectively. The aggregate offered rate is defined as $\Lambda_{\text{DL}}=\sum_{u}\lambda S_{u},$ and we report the normalized load $\rho=\Lambda_{\text{DL}}/C_{\text{DL}}$; we study both nominal load ($\rho=1.0$) and a controlled overload ($\rho=4.0$) to stress-test the scheduler’s priority protection capabilities. Class reservations (idleSlope shares) are set to {75%, 20%, 5%} for priorities $p_{1},p_{2},p_{3}$}, respectively. Per-UE credit clamps $lo_{u},hi_{u}]$ are derived following the Ethernet CBS standard [9]. The credit gate renders a UE eligible in slot $n$ iff it is backlogged, has a free HARQ process, and its credit is non-negative. HARQ retransmissions consume radio resources but do not reduce the downlink queue $Q_{u}$; therefore, only the bytes from a first transmission debit the credit. The CBS-PU mechanism relies on information readily available at the gNB. At the start of each slot, the gNB knows the queue size $Q_{u}[n]$, the chosen TBS, and the HARQ state. A new transmission service removes $\min\{\text{TBS}_{u}[n],Q_{u}[n]\}$ bytes from the buffer. Any padding added to the TB does not change $Q_{u}$.
@@ Acknowledgement
This work was funded by the Dutch ”Rijksdienst voor Ondernemend Nederland – RVO” under the TSH number 21007 – RHIADA project. The authors would also like to acknowledge the project partners, GKN Fokker Aerospace and the Netherlands Aerospace Centre (NLR), for their support and collaboration.